OpenStudy (anonymous):

Evaluate the limit as x goes to 0+ (x^2*lnx) using l'hopitals rule

OpenStudy (turingtest):

\[\lim_{x\to0^+}x^2\ln x\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I took the derivative using the product rule which was (x^2*1/x)+(2x*lnx)

OpenStudy (turingtest):

to use l'Hospital, fist we need to convert it into the form \(\frac00\) I suppose we can do the transformation \(x=\frac1t\) which implies that \(t\to\infty\) as \(x\to0^+\)

OpenStudy (turingtest):

l'hospital requires us to take the derivative of the numerator and denominator separately, not just take the derivative of the whole function like you did

OpenStudy (anonymous):

but does that mean i would need to change the form of the equation to have a denominator?

OpenStudy (turingtest):

yes, try the substitution \(x=\frac1t\) note that will change your limit though, because since we want x to approach 0, than implies that t must go to infinity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if t goes to infinity using that substitution it becomes 1 over infinity which is 0

OpenStudy (turingtest):

x goes to zero, yes and what does the whole expression become?

OpenStudy (anonymous):

0/0?

OpenStudy (turingtest):

just sub x=1/t for now no limit taking yet

OpenStudy (anonymous):

well x^2lnx=lnx/1/x^2 so using this we'll have: lim as x ->0+ lnx/1/x^2=lim as x ->0+ 1/x/(-2/x^3)=lim as x ->0+ -x^2/2=0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^+}x^2lnx=\lim_{x \rightarrow 0^+} lnx/(1/x^2)=\lim_{x \rightarrow 0^+} 1/x/(-2/x^3)\] \[=\lim_{x \rightarrow 0^+} -x^2/2=0\]

OpenStudy (turingtest):

ok, I'll give it to @anonymoustwo44 his/her way is better than mine. they both give the right answer though :)

OpenStudy (anonymous):

yep :D

OpenStudy (anonymous):

thank you both of you i just seemed to be confused with the natural log :/

OpenStudy (anonymous):

yours is unique @TuringTest

OpenStudy (turingtest):

naw, I use the \(x=\frac1t\) trick all the time so does wolfram actually :P

OpenStudy (anonymous):

does this mean that every time it is an indeterminate form the limit will be 0?

OpenStudy (turingtest):

no

OpenStudy (turingtest):

indeterminate means what it sounds like: the limit is yet to be determined

OpenStudy (anonymous):

mmmm okay

OpenStudy (anonymous):

can you guys help me with one more problem? please

OpenStudy (turingtest):

sure, could you please post separately?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is the same using l'hopitals rule limit -->1+ (x^2-1)^(x-1)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \ln(x^2)/(x^2-1)=1?\]

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

Hi, Jenjen. Welcome to Open Study =D It's considered more polite to close this question and then post a new question rather than asking in a comment.

OpenStudy (anonymous):

thanks for the welcome lol ill do that ><